OpenStudy (anonymous):

Find the rate of change for the following equation and how did you find the rate of change? 5x+3y=-2

OpenStudy (anonymous):

Differentiate it. Isolate the y's and the x's, so:\[y=\frac{ -2-5x }{ 3}\] So \[\frac{ dy }{ dx } = \frac{ d }{ dx } (\frac{ -2-5x }{ 3 })\]

OpenStudy (anonymous):

Thank you.